12:03 Powered up the DX LBTI WFS. FOund stale processes on DX WFS machine running. w_stop_all/start_all
  • Homed baysides and FW
  • Binned Ocam and cooled, binned CCD47, looks good!
  • Verified motion of camera lens
  • updated symlinks
    • Updating focus in gosetups: ln -sib LBTI_ACE-AO_20230926 LBTI_ACE-AO
    • Updating FoV for argos calunit: fov_select.py argos 17: argos (*) -35.571 -47.782
    • Verified the ace table: table_LBTI_ACE-AO.txt -> table_LBTI_ACE-AO_20230923_500m_1700Hz.txt
    • Not testing NCPA for this test so no links updated
12:08 Argos source positioned:
  • Right PSF Secondary values: -7.62, -3.93, 0.8, 242, -478, 0
  • Baysides from FoV

12:20 Close loop bin1
  • Send ACE preset 11mag
  • Although it appeared the 1.5" seeing file was uncommented for the disturbance, it appears there were multiple uncommented lines in the end. Cleaned the file up. What a mess.
  • Resent preset with 11mag and: disturbToAdd string "/local/aomeas/adsec_calib/M2C/KL_v29/atm_disturb/dist_flao2_KL_v20_atm_s1.5_L040.0_v15.0_ovfreq1700.00_sd3892.fits"
  • So even after the recent WFS update i did verify that by changing the daytime.conf the proper disturbance is loading without restarting the arb. Its only the NCPA and other symlinks that need the WFS arb to be restarted to ingest them.. the fov_select works as well, so what about text files similar to those to update?
  • Bin1 1470 Hz, closed with optical gain on. Left closed for 15 min to verify BCU47 stability
  • 193620
  • pupils well aligned with camera lens moving well. Camera lens at -48, 6. No need to redo pupils unless I really wanted to be a perfectionist, and I don't
12:40 Close loop bin2
  • Send ACE preset 12.8mag (argos source 109)
  • disturbance: disturbToAdd string "/local/aomeas/adsec_calib/M2C/KL_v29/atm_disturb/dist_flao2_KL_v20_atm_s1.5_L040.0_v15.0_ovfreq1200.00_sd3892.fits"
  • Bin2 1158Hz, closed loop with optical gain.
  • Pupils well centered: Camera lens position 13, -7
  • Left closed for an additional 14 min to monitor stability
  • 20231208_195511
12:57 Close loop bin3
  • Sending bin3 mag 14.6 (argos source 93)
  • The CCD47 closed when reconfiguring from 4x4 to 16x16, reopened without issue and looked ok
  • Disturbance: disturbToAdd string "/local/aomeas/adsec_calib/M2C/KL_v29/atm_disturb/dist_flao2_KL_v20_atm_s1.5_L040.0_v15.0_ovfreq900.00_sd3892.fits"
  • Bin3 963Hz
  • Camera lens again well centered: 3, 12
  • left closed for 10 min
  • 201015
13:12 Close loop bin4
  • Sinding preset bin4, mag 16.7 (argos source 78)
  • No issue with CCD 47 but it didn't need to rebin
  • Disturbance: disturbToAdd string "/local/aomeas/adsec_calib/M2C/KL_v29/atm_disturb/dist_flao2_KL_v20_atm_s1.5_L040.0_v15.0_ovfreq500.00_sd3892.fits"
  • OCAM went not live when reconfiguring. It was a stale display and not a real not live. Closing and reopening the GUI was all that was needed
  • When in bin3 it showed the flux int eh puils as a 16.7 but after rebinning it is showing this source as a 17.0. Calibration issue?
  • Close loop bin4 319 Hz (17.0 Mag)
  • Camera lens ok: -56, +47
  • Tip tilt grossly under estimated on first optimize gain going with near 0, better second pass. do we need to iterate more for bin4? Second pass gave TT around 0.07
  • Bin4 optimize gain first pass:
    bin4_first pass.png
  • bin4 optimize gain second pass:
    bin4 second pass.png
  • 202550, 202618 (both after second pass optimization and stable for a bit)
13:29 Last preset to bin1 to verify that CCD47 does not close after rebinning. Bin4 to bin1 successful. Bin1 to bin4 successful.

13:30 Closing up shop.
  • Argos source off
  • When hitting set safe got a blank error box. Likely because last preset was not completed but just cancelled (State AOPrepared)? I manually parked the filters and set the frame rate. This needs to be more robust. We frequently would end with a cancelled preset if in poor weather.
  • Rested DX adsec
  • Set the right PSF global offsets back to defaults
  • Reenabled lookup tables and disabled rr mode
  • w_stop_all
  • powered off 5, 8, 15, 16

-- JenniferPower - 08 Dec 2023

*
Topic revision: r1 - 08 Dec 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback